OpenStudy (anonymous):

The area of a triangular poster is 84 square feet. Its height is 2 feet shorter than the length of its base. The formula for the area of a triangle is A=1/2bh. Find the height of the triangle.

OpenStudy (jiteshmeghwal9):

let the width be x. then\[\frac{1}{2} \times (2-x) \times (x)=(84)\]

OpenStudy (anonymous):

this isn't right

OpenStudy (klimenkov):

Let the height be \(h\), then the base is \(b=h+2\). Now solve equation \(\frac 12h(h+2)=84\)

OpenStudy (klimenkov):

\(h^2+2h-168=0\)

OpenStudy (anonymous):

how did you get -168 from 84?

OpenStudy (anonymous):

I got \[h^{2}+1/2h-42\]

OpenStudy (klimenkov):

\(\frac 12h(h+2)=84\) Multiply both parts on 2: \(2\cdot\frac 12h(h+2)=2\cdot84\) \(h(h+2)=168\)

OpenStudy (anonymous):

Ok, so remove the fraction from these problems. Thanks for your help.